Find a function f (x) and a number k such that
3 + ∫_k^x▒〖f(t)dt=e^5x 〗
1
Expert's answer
2020-09-07T17:26:50-0400

Given function is "3 + \int_k^x f(t)dt = e^{5x}"


we can write it as "\int_k^x f(t)dt = e^{5x} - 3"


Differentiating both sides, we get


"f(x) = 5e^{5x}"


Now, from the equation given in question,

"3 + \int_k^x 5e^{5x}dt = e^{5x}"


We get

"3 + [e^{5x}] - e^{5k} = e^{5x} \implies k = \frac{1}{5}ln(3)"